OpenStudy (anonymous):

the ratio of the width to the height of a tv screen is 16 to 9. what is the angle between the width and a diagonal of the screen?

OpenStudy (anonymous):

tan^-1(9/16)=29.36 and tan^-1(16/9)=60.64

OpenStudy (anonymous):

90

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

well... ahemm... the pythagorean trig identities rather

OpenStudy (anonymous):

18.36

OpenStudy (jdoe0001):

 \(\bf tan(\theta)=\cfrac{16}{9}\implies tan^{-1}(\theta)=tan^{-1}\left(\cfrac{16}{9}\right)\implies \theta=tan^{-1}\left(\cfrac{16}{9}\right)\)